Solve the following problem.
The Nile River is 394 kilometers longer than the Amazon River. The two rivers have a total length of 12,986 kilometers. How long is each river?

Answer by harpazo(655)

x + x + 394 = 12,986
2x + 394 = 12, 986
2x = 12,986-394
2x = 12,592
x = 12,592/2
x = 6296
The length of one river is 6296 km and the other river is 6296 + 394 km or 6690 km.
Did you follow my notes?